An electron and a proton, having the same kinetic energy, enter a region of uniform magnetic field moving perpendicular to the field. Calculate the ratio of the radii of their circular paths in the field.
1
Expert's answer
2019-03-21T10:39:14-0400

p = 1.6x10-19 C - charge of proton

e = -1.6x10-19 C - charge of electron

me = 9.1x10-31 kg

mp = 1.6x10-27 kg

Ep=Ee

B = induction of magnetic field

Then force of Lorentz - FL=evB or for proton FL=pvB. Kinetic energy - E = mv2 / 2

mv2 / Rp = pvB, for electron mv2 / Re = evB

"\frac{R_p}{R_e} = \sqrt{\frac{m_e}{m_p}} \frac{p}{e} = 0.02384"